Circular No. 8787

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2006 Y2 (GIBBS) A. R. Gibbs reports his discovery of a comet in the course of the

Catalina Sky Survey (discovery observation tabulated below), noting that a picture from four co-added 30-s unfiltered CCD images taken with the Catalina 0.68-m Schmidt telescope on Dec. 26.5 UT shows a compact coma (stellar in appearance) with a narrow tail 12′′ long in p.a. 280^{o}; a co-added image from four 90-s co-added frames taken by Gibbs on Dec. 27.5 shows a dense coma of size 11′′ × 8′′, elongated in the direction of tail, which itself was still narrow and extending 70′′ in p.a. 290^{o}. Following posting on the Minor Planet Center’s ‘NEOCP’ webpage, M. Tichý and J. Tichá write that their CCD images taken with the 1.06-m telescope at Kleť on Dec. 27.1 and 28.1 show the comet to be diffuse. K. Sárneczky reports that a co-added image (total exposure 720 s) taken with the Konkoly Observatory 0.60-m Schmidt telescope on Dec. 27.2 shows a 16′′ tail in p.a. 280^{o}.

2006 UT  α_{2000}  δ_{2000}  Mag.
Dec. 26.48652  12^{h}46^{m}26.^{s}25  +4^{o}26′08′′.9  18.2

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2006-Y56. It is quite likely that this comet has an orbit of short period.

T = 2006 Dec. 30.971 TT  ω = 27.^{o}858
Ω = 119.437 }2000.0
q = 1.36081 AU  i = 16.188

COMETS C/2006 U12 AND C/2006 U13 (SOHO) Additional Kreutz sungrazing comets have been found on SOHO web-

site images (cf. IAUC 8786). K. Battams notes that C/2006 U12 was diffuse, small, and faint (reaching mag 7.5). C/2006 U13 was tiny, very faint, and slightly diffuse with no tail.

Comet  2006 UT  α_{2000}  δ_{2000} Inst. F  MPEC
C/2006 U12 Oct. 24.413 13^{h}49.^{m}0 −12^{o}51′ C2 TC 2006-X26
C/2006 U13  25.381 13 52.9 −13 14 C2 HS 2006-X26
2006 December 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green